Use the following information to answer questions 1-4. You have achieved a dream job working for the Toledo Mud Hens. You estQuestion 1 (0.5 points) What is the profit maximizing price? O 15 O 14 O 13 O None of the aboveQuestion 2 (0.5 points) What is the total cost at the profit maximizing price? 19000 21000 23000 None of the aboveQuestion 3 (0.5 points) What is the profit at the profit maximizing price? 76000 0 60000 47000 O None of the aboveQuestion 4 (0.5 points) Suppose some new seats are added to the stadium (ignore the cost of adding these seats for the purpos
Answer #1
Price, P     Quantity, Q TR=P*Q MR=Change in TR/Change in Q
15 5000 75000
14 6000 84000 9
13 7000 91000 7
12 8000 96000 5
11 9000 99000 3
10 10000 100000 1

1)

A monopolist will decrease the price as long as Marginal revenue of additional sale is higher or equal to marginal cost.

In this case MR ($3) > MC ($2) at Q=9000

But MR ($1) < MC ($2) at Q=10000

So, monopolist will sell 9000 tickets at a price of $11 to maximize profit (Number of tickets sold is within the capacity)

So, correct option is

None of these

2)

Total Cost=TC=5000+2*Q=5000+2*9000=$23000

Correct option is

$23000

3)

Total Revenue=TR=P*Q=11*9000=$99000

Profit=TR-TC=99000-23000=$76000

Correct option is

76000

4)

Since optimal number of tickets sold is equal to capacity of Toledo Mud Hens. Increasing capacity will not affect the optimal price.

So, correct option is

11